A circular coil of wire 0.0275 m in radius, having 25 turns, lies in a horizontal plane. It carries a current of 3.00 A in a

counterclockwise rotation when viewed from above. The coil is in the magnetic field directed towards the right with a

magnitude of 1.20 T. Find the magnetic moment and the torque on the coil.


1
Expert's answer
2022-03-20T18:48:21-0400

The magnetic moment of a coil

"\mu=NIA=NI\pi r^2\\
=25\times 3.00\times 3.14\times 0.0275^2=0.178\:\rm A\cdot m^2"

The torque

"\tau=\mu B\sin\theta\\
=0.178\times 1.20\times \sin90^\circ=0.214\: \rm N\cdot m"